OpenStudy (anonymous):

Help with following algebra question 2(p+3)-4 ANSWER 2P+2 How do they get the answer?

OpenStudy (anonymous):

dISTRIBUTE

OpenStudy (anonymous):

2(p+3)-4 Start with the parenthesis 2p+6-4= 2p+2

OpenStudy (anonymous):

2P+6-4 = 2P+ 2 = 2(P+1)

OpenStudy (anonymous):

Where does the 6 come from? Sorry but you have to show me working out?

OpenStudy (anonymous):

2×3=6

OpenStudy (aripotta):